A ball is thrown vertically upward from the ground with a velocity of 30 m/s. When is the displacement of the ball zero?


1
Expert's answer
2021-07-29T01:19:07-0400

The displacement of the ball would be zero, when it will be at the ground:

"t=2*\frac{v_0}{g}=2*\frac{30\:\rm m/s}{9.8\:\rm m/s^2}=6.1\:\rm s"